• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:34:13 on October 19, 1998

    Entry number 3154

    keyword=Beam Time Accounting
    Hall A

    10/18/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: Pibero Djawotho
    Shift Workers: Bill Hersman, Tim Black

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 0 0 0
    Beam Available
    but Not in Use
    1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 8
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    0 0 0 0 0 0 0 0 0
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.08
    Hall not
    ready reasons
    Target 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 8
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: Recovery from target burst is underway. Q-curve of pick-up coils. Water calibration. Cosmic run.

    Data Entry by: Pibero Djawotho